OpenStudy (anonymous):

Which equation represents a nonlinear function? A. y = x + 7 B. y = 3x – 1 C. y = 3x^2 + 1 D. y = x – 1/2

OpenStudy (zzr0ck3r):

technally all of them are non linear, linear must contain the origin. The question shuold have asked which one is not affine. correct your teacher:)

OpenStudy (anonymous):

@zzr0ck3r How dumb are you linear functions do not have to contain the origin I want to meet your teacher Y-INTERCEPT DUMB Y=MX+B

OpenStudy (zzr0ck3r):

they do, its a common misconception...look it up. b must = 0 for them to be affine. ie f(x+y) = f(x) + f(y)

OpenStudy (zzr0ck3r):

http://mathforum.org/kb/message.jspa?messageID=242695 feel free to read befor you call people dumb...

OpenStudy (anonymous):

This is BASIC ALGEBRA

OpenStudy (zzr0ck3r):

I meant to say b must equal 0 for it to be linear...

OpenStudy (anonymous):

NOT PHYSICS

OpenStudy (zzr0ck3r):

Again dont tell me im wrong, go look it up...

OpenStudy (anonymous):

I know that in physics it is but it is irrelevant to say that at this point

OpenStudy (zzr0ck3r):

no its not, its math.....by definition ....

OpenStudy (anonymous):

But the closest answer is C, right?

OpenStudy (zzr0ck3r):

most of math is the study of vector spaces...all of math....so its a must.

OpenStudy (anonymous):

-________________________________________________________________________- "In the fields of physics and engineering, which is where I'm coming from,"

OpenStudy (zzr0ck3r):

correct the intended answer is C

OpenStudy (zzr0ck3r):

again I dont care about what someone posted on a site. look up the deffinition. http://en.wikipedia.org/wiki/Linear_function

OpenStudy (anonymous):

Alright, thanks guys

OpenStudy (anonymous):

"So, all affine combinations are linear combinations but not conversely."

OpenStudy (anonymous):

FROM YOUR SOURCE ~.~

OpenStudy (zzr0ck3r):

In mathematics, a linear function means a function that is a linear map, that is, a map between two vector spaces that preserves vector addition and scalar multiplication. For example, if and are represented as coordinate vectors, then the linear functions are those functions that can be expressed as where M is a matrix. A function is a linear map if and only if = 0. For other values of this falls in the more general class of affine maps.

OpenStudy (zzr0ck3r):

here m is a 1x1 matrix

OpenStudy (zzr0ck3r):

point being just because someone on a site says this is why there is a common mmisconception does not mean we should ignore it.

OpenStudy (zzr0ck3r):

the same page you took that from explains the definition, so dont pick the quote that says nothing is my point

OpenStudy (zzr0ck3r):

anyway ...now you know:)